# WI2easy: warm inflation dynamics made easy

### v1

@@ 3 The perturbation equations in WI
where $\Gamma(x)$ is the Gamma-function, the subindex * means that all quantities are to be evaluated at the Hubble crossing time, e.g., at the moment where the comoving mode satisfies $k=aH$, and $n_{*}$ accounts for the possibility of a thermal distribution of the inflaton field due to the presence of the radiation bath. When the inflaton perturbations are fully thermalized, $n_{*}$ assumes the Bose-Einstein distribution, $n_{*}\equiv n_{\rm BE}=1/[\exp(H/T)-1]$ (for intermediate cases, see, e.g. ref. [42]).
To overcome the above mentioned pitfalls in the derivation of eq. (3.1), the common practice [36, 37, 38] is to solve the complete set of coupled perturbation equations and correct the above analytical result by a function $G(Q)$, derived from the numerical solution of the full coupled perturbation equations. The function $G(Q)$ is then determined through a numerical fitting, obtained by comparing the numerical power spectrum with the analytical proxy one, eq. (3.1),
@@
A second point of relevance in our code implementation is the inclusion or not of the term $n_{*}$ (which we set as in the fully thermalized case, $n_{*}=n_{\rm BE}$) in the power spectrum expression eq. (3.1). Which one of these two possibilities would apply, it depends on the details of the microphysics of the problem under study and whether scattering rates of the inflaton with other fields happen at a rate $\Gamma_{\rm scat}$ fast enough compared to the expansion rate, $\Gamma_{\rm scat}>H$ to ensure thermalization. In most applications in WI, where such analysis is not available, it is typically considered both possibilities (non thermalized inflaton perturbations, $n_{*}=0$, or fully thermalized, $n_{*}=n_{\rm BE}$) and, hence, the results are presented when considering each one of them. In the absence of the thermal distribution term, we can simply solve the inflaton perturbation equation assuming standard Bunch-Davis initial conditions. This ensures that the homogeneous part of the solution for $\delta\phi$ will reproduce the CI result $\langle|\delta\phi|\rangle=H/(2\pi)$. When considering the presence of the thermal contribution $n_{\rm BE}$, we make use of the proposal of ref. [30], of changing the perturbation equation for $\delta\phi$, eq. (LABEL:deltaddotphi) and adding to the right-hand-side of it another Gaussian stochastic noise term [^3] Note that in this case, we can just set the initial conditions for the inflaton perturbations to zero, as the extra noise term already ensures that the correct CI result for the spectrum spectrum appears from the particular solution as a consequence of the extra noise term.,
@@
where the noise term $\xi^{(q)}$ satisfies a two-point correlation function like (3.18) and the amplitude of the noise in (3.19) is such that to correctly reproduce the form of the power spectrum in eq. (3.1). Here, we avoid the prescription considered in ref. [41] of multiplying the part of the spectrum coming from the homogeneous solution for $\delta\phi$ by a factor $(1+2n_{*})$ at end. The prescription adopted here concerning this thermal contribution to the power spectrum is also the same one considered in the WarmSPy code (see ref. [39] for details of their implementation). In our code, we also make available the option of including or not the thermal term $n_{*}$ in the numerical simulations.
@@
[…]
@@ 4.2.1 Command ParametersWI[]
    Potential shape: specify which type of model you are working with, e.g. large field, small field or hybrid model (in the example of the quartic potential, it is a large field potential)555 The classification of large or small field potentials is the usual from the literature. Any potential where the inflaton starts at a large value and runs down to a minimum at $\phi=0$ is a large field model (e.g. monomial chaotic potentials, Starobinsky potential, $\alpha$-attractor potentials in general, all fall in this category). Otherwise, if the inflaton field starts in a plateau around the origin ($\phi=0$) and runs towards large field values, it is a small field potential (e.g. hilltop potentials, axion type of potentials, etc). If the choice is a hybrid potential, then specify the critical value for the inflaton field, which determines where inflation will end (do not add the waterfall field. This is still considered a one-field model).
@@ 4.2.2 Command FindICs[Qi,Qf]
By running the command FindICs[Qi,Qf], the background equations are solved and the code will run the search algorithm to find the appropriate initial conditions (ICs) for the model and that lead to the specified total duration in e-folds. The ICs will be generated for values of the dissipation ratio in the interval $[Q_{i},Q_{f}]$. The results will be saved in the data file starting with ICs__\__data. Here, another data file is also produced and saved with the name cp__\__data, which contains the values of $c$ and $p$ defined by eqs. (3.16) and (3.17), respectively. The results for $c$ and $p$ are used in subsequent parts of the code.
@@ 4.2.3 Command FindGQ[]
After generating the ICs in the previous stage, it can now be issued the command FindGQ[]. Here, the code will make use of the ICs in the interval $[Q_{i},Q_{f}]$ set in the previous command, run the perturbation equations using eq. (3.31) and compute the power spectrum from eq. (3.28). For each value of $Q\in[Q_{i},Q_{f}]$, it is used a value of scale $k$ in the perturbation equations that corresponds to seven e-folds before Hubble radius crossing (whose values are also computed when running the previous command FindICs[Qi,Qf]), which correspond to a value around $k\simeq 10^{3}a(t_{i})H(t_{i})$). The perturbations are then evolved an additional seven e-efolds, which ensures the spectrum has already been frozen and $k\ll aH$. The obtained numerical power spectrum is then compared to the analytical proxy result eq. (3.1) and the function $G(Q)$ computed from the definition eq. (3.2). The results are saved in the data file starting with G(Q)__\__data. Note that the power spectrum that is evaluated at this stage is still the non-normalized one.
@@ 4.2.4 Command ObservationsWI[Q0value,x0value]
ObservationsWI[Q0value,x0value] will execute the following steps: (1) for the value of Q0value for $Q$ and starting from the initial value x0value for the inflaton, the code will search for the appropriate conditions leading to the total number of e-efolds set at the beginning of the code, but this time it will also make sure that the power spectrum at the initial number of e-folds will have the correct amplitude. This way, the correct value $V_{0}$ for the normalization of inflaton potential is determined. Here, the code makes uses of the proxy power spectrum eq. (3.1) but corrected by the function $G(Q)$. Instead of using any numerical fitting for the data for $G(Q)$ generated from the previous command FindGQ[], the code makes use of the raw data from the G(Q)__\__data file and does a spline interpolation of the data to create a function $G(Q)$ that can be used internally in the calculations. At this stage, the code still makes use of the slow-roll conditions and the approach described at the end of sec. 2. The code performs then the second step. (2) To avoid any issues of numerical precision due to the use of the approximated equations (and e.g. the field trajectory not to be in the inflationary attractor one), the code will run the background equations backwards for about seven e-efolds and run it again forward from that point until the end of inflation. The point $N_{*}$ where the pivot scale crosses the Hubble radius (where the amplitude of the power spectrum at the pivot scale is given by the value $A_{s}$), the point where inflation ends $N_{\rm end}$ (i.e., where $\epsilon_{H}=1$) and also the point $N_{\rm reh}$ where the universe transits to the radiation dominated regime after inflation (i.e., where $\epsilon_{H}\simeq 2$) are determined. The code also verifies whether the value of $N_{*}$ satisfies the correct point where the pivot scale have crossed the Hubble radius, given by the condition [24]
@@ 4.2.5 Command QrangeWI[Qinit]
The code command QrangeWI[Qinit] will repeat the execution of the previous command, varying $Q$ from an initial value $Q_{i}$ up to the largest value of $Q$ available from the data file ICs__\__data generated by the FindICs command. The command will stop in case $n_{s}$, for some value of $Q$, gets outside of the range $n_{s}\in[0.93,0.99]$. Results for the background and perturbation quantities at the Hubble radius crossing $N_{*}$ for each value of $Q$ considered are then saved in the data file observables__\__data. Examples of several plots that can be made with these results are given in the last section of the companion Mathematica notebook main.nb.
@@ 5 Models analyzed
    $V(\phi,\sigma)=\frac{1}{4\lambda}(M^{2}-\lambda\sigma)^{2}+\frac{1}{2}m^{2}\phi^{2}+\frac{1}{2}g^{2}\phi^{2}\sigma^{2}.$
@@ 6.2 Comparison with previous results for $G(Q)$
In the fig. 5 we show the comparison of our results with these different earlier forms of fitting functions for $G(Q)$ found in the literature. We focus on the two most well used fitting functions for $G(Q)$ used for the dissipation coefficients $\Upsilon\propto T$ and $\propto T^{3}$ (see, e.g. refs. [11, 19]) and also the fittings found from WarmSPy, ref. [39], for these same cases. The earlier fitting functions tend to under estimate [^9] We note that the earlier fitting formulas for $G(Q)$ were designed to better fit the power spectrum in the low $Q$ region. The reason for this was that the existing models of WI at the time were only able to produce results consistent with the observations in the weak dissipative regime [11]. $G(Q)$ for $Q\gtrsim 50$. The most recent fittings from ref. [39] provide a better coverage in a much higher range of values for $Q$.
As already commented before, here we do not offer any new fitting formulas for $G(Q)$. WI2easy, through the command FindGQ[] generates sufficient data points from which an internal function $G(Q)$ is created through a spline interpolation. This is then used in the subsequent parts of the code, e.g. in the modules ObservationsWI and QrangeWI. We find this approach to be more efficient and to lead to better precision when obtaining the perturbation quantities that are direct compared with the observation data.
@@ 6.4 Dependence of the background and perturbation quantities on the value of $Q$
Figure 9: Perturbation quantities when varying $Q_{*}$. The top gray region in panel (a) and the left gray region in panel (b) indicate the region where $r>0.036$, with the upper bound on $r$ from the BICEP, Keck Array and Planck combined data [63]. The cyan regions in panels (b) and (c) indicate the one- and two-sigma constrains also from ref. [63].
Finally, in fig. 9 we give the perturbation quantities obtained for the quartic monomial potential with a linear in the temperature dissipation coefficient (and considering thermal inflaton perturbations and the radiation noise term in the radiation perturbation equation). The cyan regions in the figure are the one- and two-sigma constraints [^10] Note that recently, the Atacama Cosmology Telescope (ACT) have reported new constraints which shifts the spectral tilt towards slight larger values. In WI, and in special in the example model studied here, this implies that slight larger values of $Q_{*}$ are allowed, increasing the viability of WI towards the strong dissipation regime [64]. obtained from the BICEP, Keck Array and Planck combined data [63].
@@ 7 Conclusion
The code advances the analysis of WI perturbations by employing a Fokker-Planck formalism, bypassing the direct numerical solution of stochastic differential equations. This approach transforms the computation of the primordial power spectrum–a statistical quantity derived from two-point correlation functions–into the deterministic evolution of a coupled matrix differential equation system, integrated alongside the background equations. Compared to traditional stochastic methods for WI perturbations, this framework offers significant computational advantages: (a) there is a gain in speed – the deterministic system eliminates the need for ensemble averaging over stochastic realizations and (b) precision – leveraging Wolfram Mathematica’s high-precision numerical solvers, the code allows users to tailor integration methods (e.g., adaptive step sizes, stiffness handling) to their specific model requirements. Additionally, the code prioritizes accessibility. Its modular design ensures straightforward execution for standard use cases while retaining flexibility–users can easily modify default parameters or implement model-specific adjustments without advanced programming expertise.
@@ Acknowledgments
R.O.R. acknowledges financial support by research grants from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grant No. 307286/2021-5, and from Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/201.150/2021. The work of G.S.R. is supported by a PhD scholarship from FAPERJ.

### v2

@@ 3 The perturbation equations in WI
where $\Gamma(x)$ is the Gamma-function, the subindex * means that all quantities are to be evaluated at the Hubble crossing time, e.g., at the moment where the comoving mode satisfies $k=aH$, and $n_{*}$ accounts for the possibility of a thermal distribution of the inflaton field due to the presence of the radiation bath. When the inflaton perturbations are fully thermalized, $n_{*}$ assumes the Bose-Einstein distribution, $n_{*}\equiv n_{\rm BE}=1/[\exp(H/T)-1]$ (for intermediate cases, see, e.g. ref. [42]). Here we must note that the presence or not of this thermal contribution term is controversial in the literature. In principle, one would expect that its presence or not should be strongly dependent on the details of the microphysics of the specific WI model under study, which would involve, for example, whether scattering rates of the inflaton with other fields happen at a rate $\Gamma_{\rm scat}$ fast enough compared to the expansion rate, $\Gamma_{\rm scat}>H$, such as to ensure thermalization of the inflaton [^3] We note that [16] has also presented arguments about the absence of this thermal contribution to the power spectrum, even when the inflaton thermalizes. However, the presence of the thermal bath in WI might give origin to other type of corrections to the scalar and tensor spectra that are of order ${\cal O}(H^{3}/T^{3})$, as shown in [43].. We believe that the presence or not of this term can only be settled after an appropriate lattice simulation of the full dynamics, or dedicated kinetic Botzmann equation study for the occupation number for the inflaton in WI is performed. While this dedicated analysis is still lacking, most applications in WI typically considered both possibilities (non thermalized inflaton perturbations, $n_{*}=0$, or fully thermalized, $n_{*}=n_{\rm BE}$) and, hence, the results are presented when considering each one of them. In our code, we give the option to include or not this term and users of the code can decide what would be more appropriate for their analysis.
While we do not expect eq. (3.1) to match the full numerical result for the power spectrum in WI, we can overcome this shortcoming by using a common practice [36, 37, 38], which is to solve the complete set of coupled perturbation equations and correct the analytical expressions eq. (3.1) by a function $G(Q)$, derived from the numerical solution of the full coupled perturbation equations. The function $G(Q)$ is then determined through a numerical fitting, obtained by comparing the numerical power spectrum with the analytical proxy one, eq. (3.1),
@@
A second point of relevance in our code implementation is the inclusion or not of the term $n_{*}$ (which we set as in the fully thermalized case, $n_{*}=n_{\rm BE}$) in the power spectrum expression eq. (3.1). As already mentioned, the presence or not of this terms is strongly dependent on the details of the microphysics of the specific WI model under study. In the absence of the thermal distribution term, we can simply solve the inflaton perturbation equation assuming standard Bunch-Davis initial conditions. This ensures that the homogeneous part of the solution for $\delta\phi$ will reproduce the CI result $\langle|\delta\phi|\rangle=H/(2\pi)$. When considering the presence of the thermal contribution $n_{\rm BE}$, we make use of the proposal of ref. [30], of changing the perturbation equation for $\delta\phi$, eq. (LABEL:deltaddotphi) and adding to the right-hand-side of it another Gaussian stochastic noise term [^4] Note that in this case, we can just set the initial conditions for the inflaton perturbations to zero, as the extra noise term already ensures that the correct CI result for the spectrum spectrum appears from the particular solution as a consequence of the extra noise term.,
@@
where the noise term $\xi^{(q)}$ satisfies a two-point correlation function like (3.18) and the amplitude of the noise in (3.19) is such that to correctly reproduce the form of the power spectrum in eq. (3.1). Here, we avoid the prescription considered in ref. [41] of multiplying the part of the spectrum coming from the homogeneous solution for $\delta\phi$ by a factor $(1+2n_{*})$ at the end. The prescription adopted here concerning this thermal contribution to the power spectrum is also the same one considered in the WarmSPy code (see ref. [39] for details of their implementation). In our code, we also make available the option of including or not the thermal term $n_{*}$ in the numerical simulations.
@@
where $\langle\ldots\rangle$ means here the ensemble average over the different realization of the noise terms in the perturbation equations and $\mathcal{R}$ is defined as (in the uniform-curvature gauge)
@@ 4.2.1 Command ParametersWI[]
    Potential shape: specify which type of model you are working with, e.g. large field, small field or hybrid model (in the example of the quartic potential, it is a large field potential)666 The classification of large or small field potentials is the usual from the literature. Any potential where the inflaton starts at a large value and runs down to a minimum at $\phi=0$ is a large field model (e.g. monomial chaotic potentials, Starobinsky potential, $\alpha$-attractor potentials in general, all fall in this category). Otherwise, if the inflaton field starts in a plateau around the origin ($\phi=0$) and runs towards large field values, it is a small field potential (e.g. hilltop potentials, axion type of potentials, etc). If the choice is a hybrid potential, then specify the critical value for the inflaton field, which determines where inflation will end (do not add the waterfall field. This is still considered a one-field model).
@@ 4.2.2 Command FindICs[Qi,Qf]
By running the command FindICs[Qi,Qf], the background equations are solved and the code will run the search algorithm to find the appropriate initial conditions (ICs) for the model and that lead to the specified total duration in e-folds. The ICs will be generated for values of the dissipation ratio in the interval $Q_{i},Q_{f}]$. The results will be saved in the data file starting with ICs_\_data. Here, another data file is also produced and saved with the name cp_\_data, which contains the values of $c$ and $p$ defined by eqs. (3.16) and (3.17), respectively. The results for $c$ and $p$ are used in subsequent parts of the code.
@@ 4.2.3 Command FindGQ[]
After generating the ICs in the previous stage, it can now be issued the command FindGQ[]. Here, the code will make use of the ICs in the interval $Q_{i},Q_{f}]$ set in the previous command, run the perturbation equations using eq. (3.31) and compute the power spectrum from eq. (3.28). For each value of $Q\in[Q_{i},Q_{f}]$, it is used a value of scale $k$ in the perturbation equations that corresponds to seven e-folds before Hubble radius crossing (whose values are also computed when running the previous command FindICs[Qi,Qf]), which correspond to a value around $k\simeq 10^{3}a(t_{i})H(t_{i})$). The perturbations are then evolved an additional seven e-efolds, which ensures the spectrum has already been frozen and $k\ll aH$. The obtained numerical power spectrum is then compared to the analytical proxy result eq. (3.1) and the function $G(Q)$ computed from the definition eq. (3.2). The results are saved in the data file starting with G(Q)_\_data. Note that the power spectrum that is evaluated at this stage is still the non-normalized one.
@@ 4.2.4 Command ObservationsWI[Q0value,x0value]
ObservationsWI[Q0value,x0value] will execute the following steps: (1) for the value of Q0value for $Q$ and starting from the initial value x0value for the inflaton, the code will search for the appropriate conditions leading to the total number of e-efolds set at the beginning of the code, but this time it will also make sure that the power spectrum at the initial number of e-folds will have the correct amplitude. This way, the correct value $V_{0}$ for the normalization of inflaton potential is determined. Here, the code makes uses of the proxy power spectrum eq. (3.1) but corrected by the function $G(Q)$. Instead of using any numerical fitting for the data for $G(Q)$ generated from the previous command FindGQ[], the code makes use of the raw data from the G(Q)_\_data file and does a spline interpolation of the data to create a function $G(Q)$ that can be used internally in the calculations. At this stage, the code still makes use of the slow-roll conditions and the approach described at the end of sec. 2. The code performs then the second step. (2) To avoid any issues of numerical precision due to the use of the approximated equations (and e.g. the field trajectory not to be in the inflationary attractor one), the code will run the background equations backwards for about seven e-efolds and run it again forward from that point until the end of inflation. The point $N_{*}$ where the pivot scale crosses the Hubble radius (where the amplitude of the power spectrum at the pivot scale is given by the value $A_{s}$), the point where inflation ends $N_{\rm end}$ (i.e., where $\epsilon_{H}=1$) and also the point $N_{\rm reh}$ where the universe transits to the radiation dominated regime after inflation (i.e., where $\epsilon_{H}\simeq 2$) are determined. The code also verifies whether the value of $N_{*}$ satisfies the correct point where the pivot scale have crossed the Hubble radius, given by the condition [24] [^7] In practice, the code always check whether eq. (4.1) is satisfied or not, adjusting $N_{tot}$ and the initial conditions until the condition given by eq. (4.1) is fulfilled. Although most of the evolution for the background quantities is presented starting from the point $N_{*}$ (which follows the usual convention of always meaning the number of e-folds before the end of inflation for which the pivot scale crosses the Hubble radius), it is also useful to present results before this point (e.g. for $N_{tot}>N_{*}$), like, for example, for the evolution of the power spectrum (see, e.g. fig. 7(a) below). This way we can cover scales both below and above the pivot scale.
@@ 4.2.5 Command QrangeWI[Qinit]
The code command QrangeWI[Qinit] will repeat the execution of the previous command, varying $Q$ from an initial value $Q_{i}$ up to the largest value of $Q$ available from the data file ICs_\_data generated by the FindICs command. The command will stop in case $n_{s}$, for some value of $Q$, gets outside of the range $n_{s}\in[0.93,0.99]$. Results for the background and perturbation quantities at the Hubble radius crossing $N_{*}$ for each value of $Q$ considered are then saved in the data file observables_\_data. Examples of several plots that can be made with these results are given in the last section of the companion Mathematica notebook main.nb.
@@ 5 Models analyzed
    $V(\phi,\sigma)=\frac{1}{4\lambda}(M^{2}-\lambda\sigma^{2})^{2}+\frac{1}{2}m^{2}\phi^{2}+\frac{1}{2}g^{2}\phi^{2}\sigma^{2}.$
@@ 6.2 Comparison with previous results for $G(Q)$
In Fig. 5, we compare our results with earlier parameterizations of $G(Q)$ from the literature, focusing on the two most widely used forms for dissipation coefficients $\Upsilon\propto T$ and $\propto T^{3}$ (see, e.g. refs. [11, 19]) as well as recent fits derived from WarmSPy [39]. The earlier fitting functions tend to under estimate [^11] We note that the earlier fitting formulas for $G(Q)$ were designed to better fit the power spectrum in the low $Q$ region. The reason for this was that the existing models of WI at the time were only able to produce results consistent with the observations in the weak dissipative regime [11]. $G(Q)$ for $Q\gtrsim 50$. The most recent fittings from ref. [39] accommodate recent WI scenarios where strong dissipation enhances observational signatures (e.g., suppressed tensor modes). Our analysis highlights the necessity of updated $G(Q)$ parameterizations for precision cosmology in high-$Q$ regimes, particularly for temperature-sensitive dissipation coefficients.
As already commented before, here we do not offer any new fitting formulas for $G(Q)$. WI2easy, through the command FindGQ[], generates sufficient data points from which an internal function $G(Q)$ is created through a spline interpolation. This is then used in the subsequent parts of the code, e.g. in the modules ObservationsWI and QrangeWI. We find this approach to be more efficient and to lead to better precision when obtaining the perturbation quantities that can be direct compared with the observation data.
@@ 6.4 Dependence of the background and perturbation quantities on the value of $Q$
Figure 9: Perturbation quantities when varying $Q_{*}$. The top gray region in panel (a) and the left gray region in panel (b) indicate the region where $r>0.036$, with the upper bound on $r$ from the BICEP, Keck Array and Planck combined data [64]. The cyan regions in panels (b) and (c) indicate the one- and two-sigma constraints, also from ref. [64].
Finally, in fig. 9 we give the perturbation quantities obtained for the quartic monomial potential with a linear in the temperature dissipation coefficient (and considering thermal inflaton perturbations and the radiation noise term in the radiation perturbation equation) [^12] This result also compares well with the one obtained in [41], shown by the purple line in fig. 11 of that reference, except that here we have chosen $g_{*}=100$, while in that reference the authors considered $g_{*}=12$, as motivated from [12].. The cyan regions in the figure are the one- and two-sigma constraints [^13] Note that recently, the Atacama Cosmology Telescope (ACT) have reported new constraints which shifts the spectral tilt towards slight larger values. In WI, and in special in the example model studied here, this implies that slight larger values of $Q_{*}$ are allowed, increasing the viability of WI towards the strong dissipation regime [65]. obtained from the BICEP, Keck Array and Planck combined data [64].
@@ 7 Conclusion
The code advances the analysis of WI perturbations by employing a Fokker-Planck formalism (originally formulated in [40, 41]), bypassing the direct numerical solution of stochastic differential equations. This approach transforms the computation of the primordial power spectrum–a statistical quantity derived from two-point correlation functions–into the deterministic evolution of a coupled matrix differential equation system, integrated alongside the background equations. Compared to traditional stochastic methods for WI perturbations, this framework offers significant computational advantages: (a) there is a gain in speed – the deterministic system eliminates the need for ensemble averaging over stochastic realizations and (b) precision – leveraging Wolfram Mathematica’s high-precision numerical solvers, the code allows users to tailor integration methods (e.g., adaptive step sizes, stiffness handling) to their specific model requirements. Additionally, the code prioritizes accessibility. Its modular design ensures straightforward execution for standard use cases while retaining flexibility–users can easily modify default parameters or implement model-specific adjustments without advanced programming expertise.
@@ Acknowledgments
R.O.R. acknowledges financial support by research grants from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grant No. 307286/2021-5, and from Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/201.150/2021. The work of G.S.R. is supported by a PhD scholarship from FAPERJ. The authors acknowledge Wolfram for providing a Mathematica for sites software license.